OpenStudy (anonymous):

Mr. Brown bought a television set and paid $110 as the first installment. If he increases each payment by $20 and pays a total of 6 installments, how much will the television set have cost him?

OpenStudy (anonymous):

It would be in the form of 110 + 20n, because with each installment the payment will increase an additional 20 dollars. There is a total of 6, so the sum of each installment needs to be added. So take the sum for each n up to 6 should be right!

OpenStudy (texaschic101):

110 for first installment. (1st installment) 110 + 20 = 130 (2nd installment) 130 + 20 = 150 (3rd installment) 150 + 20 = 170 (4th installment) 170 + 20 = 190 (5th installment) 190 + 20 = 210 (6th installment) I got 210. There is probably a better way to do this...I just don't know it.